# Beyond the Labels: Unveiling Text-Dependency in Paralinguistic Speech Recognition Datasets

### v1

@@ Abstract
Paralinguistic traits like cognitive load and emotion are increasingly recognized as pivotal areas in speech recognition research, often examined through specialized datasets like CLSE and IEMOCAP. However, the integrity of these datasets is seldom scrutinized for text-dependency. This paper critically evaluates the prevalent assumption that machine learning models trained on such datasets genuinely learn to identify paralinguistic traits, rather than merely capturing lexical features. By examining the lexical overlap in these datasets and testing the performance of machine learning models, we expose significant text-dependency in trait-labeling. Our results suggest that some machine learning models, especially large pre-trained models like HuBERT, might inadvertently focus on lexical characteristics rather than the intended paralinguistic features. The study serves as a call to action for the research community to reevaluate the reliability of existing datasets and methodologies, ensuring that machine learning models genuinely learn what they are designed to recognize.
@@ I Introduction
While the primary focus of speech recognition research gravitates towards Automatic Speech Recognition (ASR), the study of paralinguistic traits, such as cognitive load, physiological stress and emotions, remains a significant field too. These traits are of interest for applications, ranging from human-computer interaction to psychological research, and rely heavily on dedicated datasets. One such dataset is Cognitive Load with Speech and EGG (CLSE[1], commonly used for cognitive load recognition. Another key dataset is the  Interactive Emotional Dyadic Motion Capture (IEMOCAP[2], primarily employed for emotion recognition.
A prevailing assumption is that machine learning algorithms trained on these datasets learn to recognize paralinguistic traits based on observable physiological or psychological changes in speech production.
However, this paper challenges this assumption by revealing a critical oversight: we provide evidence of significant lexical correlation between the labels (e.g., cognitive load or emotion) and the uttered sentences within these datasets.
We extensively analyse CLSE and review the design of IEMOCAP to substantiate this claim. Given the implications of our findings, the paper serves as a call to action for the research community. It urges a reevaluation of existing datasets and methodologies to ascertain that machine learning systems are learning to recognize what they are designed to recognize.
II Background
@@ II-A Datasets in Focus
  • CLSE (Cognitive Load with Speech and EGG): provides an approach to recognizing cognitive load by incorporating both speech recordings and Electroglottograph (EGG) signals.
@@
  • IEMOCAP (Interactive Emotional Dyadic Motion Capture) [^1] https://sail.usc.edu/iemocap/index.html: covers about twelve hours of scripted and spontaneous dialogues. It captures detailed facial and hand movements to study a range of emotions.
II-B Machine Learning Approaches
@@
Various machine learning algorithms, from classical UBM-iVector [3], through LSTM [4] to recent large pre-trained models like wav2vec [5], have been employed on these datasets. The general presumption is that these models are capturing paralinguistic features rooted in physiological or psychological changes [6, 7].
II-C The What and the How
@@
It is crucial to note that in paralinguistic research, both the content (what) and manner (how) of speech carry significance. While the physiological and prosodic features often capture the how, the lexical choices, semantics, and syntax reveal insights into the what. Ignoring either aspect could result in an incomplete understanding of paralinguistic traits. Therefore, this study highlights the importance of acknowledging the lexical correlates when analyzing paralinguistic datasets.
III CLSE: Methodology
III-A Dataset and Task Description
@@
Our primary study focuses on the main part of the CLSE database, termed as CLSE-Span, which employs the  Reading Span task (RSPAN[8] task. In RSPAN, participants are asked to validate sentences for their logical coherence while memorizing a sequence of letters. Sentences are randomly chosen from a closed set. The task is divided into sets, where each set contains varying trials of sentence validation and letter memorization. The database consists of 21 such sets with each participant producing 75 utterances, totaling 1800 utterances. The average utterance duration is 4 seconds. Labels for the cognitive load values are: low, medium and high load.
The metric for this study is the Unweighted Average Recall (UAR). UAR offers a balanced assessment of performance by averaging recall across all classes, making it particularly valuable for evaluating imbalanced datasets.
III-B Experimental setup
To explore the dataset, we initially employed ASR using Whisper [9] with a pre-trained multilingual large-v3 model to transcribe the utterances. Subsequently, we utilized Sentence Transformer embeddings as features. We used the Language-Agnostic BERT Sentence Embeddings (LaBSE) model [10].
@@ III-B Experimental setup
LaBSE utilizes a dual-encoder architecture trained in two steps: first on 17 billion monolingual sentences via Masked Language Modeling, and then further refined on 6 billion translation pairs covering 109 languages. The final model is publicly accessible [^2] https://huggingface.co/sentence-transformers/LaBSE and was used to extract 768-dimensional sentence embeddings.
We then applied Agglomerative Hierarchical Clustering (AHC) on the generated utterance embeddings. After manual correction of minor clustering errors, we have identified 81 different sentences being spoken. Ground truth sentences are not available in the distributed CLSE version. The cluster size ranged between 12 and 26 utterances with majority size of 26 utterances. Given the cluster membership, we have assigned sentence ID to each utterance.
@@
For each cluster, we identified the most frequent cognitive load label in the training set. These identified labels were then used as a proxy to predict the cognitive load of utterances in the test and validation sets based on their cluster identity, effectively bypassing the use of any machine learning model for classification. This approach revealed how well the dataset’s original labels correlated with the cluster-based predictions.
@@
IV CLSE:Results
IV-A Cluster-Based Classification: Impact of Lexical Overlap
We assessed the influence of lexical overlap on classification performance using both original and fixed (shuffled) CLSE data splits. The original CLSE data displayed good performance metrics, unfortunately revealing pattern in the validation set that confirmed lexical factors’ role.
@@ IV-A Cluster-Based Classification: Impact of Lexical Overlap
After minimizing lexical redundancy in shuffled splits, the performance metrics notably decreased down to the chance level, as can be seen in Table I, further substantiating the lexical overlap’s impact on classification.
IV-B UBM-ivector system: ComPaRE 2014 replication
We have reached out to the winners of the ComPaRE 2014 [11] challenge and they helpfully provided us with the system setup of their winning system. Using their code, we were able to replicate and further improve on their results.
The UBM-iVector system uses a 64-component Universal Background Model (UBM) with 50 dimensional iVector extractor and Support Vector Machine (SVM) for classification. The optimal size for the system components were found by running grid search for UBM, iVector Extractor and SVM parameters. The details of the original system architecture can be found in [3].
Using the same approach as in Section IV-A, we have obtained results which shows significant degradation of the performance on validation and test set, see Table II.
@@ IV-B UBM-ivector system: ComPaRE 2014 replication
IV-C HuBERT based system
To corroborate our findings with a more recent speech-based system, we used HuBERT [12] with pre-trained large-ll60k weights trained on 60k hours from the Libri-Light dataset [13] as a feature extractor. As a classifier, we used Attentive correlation pooling layer from [14]. The results of our experiments are shown in Table III.
@@ IV-C HuBERT based system
The results show the same trend as seen in the text-based and UBM-iVector based systems. Accros the board, removing the lexical context hurts the performance with at least absolute 12% UAR degradation.
V IEMOCAP: Methodology
V-A Dataset and Task Description
As a secondary experiment, we have analyzed IEMOCAP dataset. It comprises 151 dialogue videos, featuring a pair of speakers in each session, resulting in a total of 302 individual videos with aproximately 12 hours of data. Ten actors were recorded in dyadic sessions (5 sessions with 2 subjects each). They were asked to perform three selected scripts with clear emotional content. In addition to the scripts, the subjects were also asked to improvise dialogs in hypothetical scenarios, designed to elicit specific emotions (happiness, anger, sadness, frustration and neutral state).
Annotations were expanded against the original set of classes, to include the identification of nine distinct emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral), along with metrics for valence, arousal, and dominance.
V-B Experimental setup
As the structure of the database is different from CLSE, we have taken slightly modified approach to verify our hypothesis. The recording sessions were either improvised or scripted, with 12 different improvised scenarios and 3 scripts. We argue that the type of scenario in the recording session represents enough lexical context to the classifier that it can diminish the other factors (prosody etc.).
To verify such claim, we concieved three different experiments. In the first one, we computed the most probable label per given session in a similar fashion to III-B. Then we obtained ASR transcripts with Whisper ASR (large-v3) and incorporated the sentence embeddings using the same extractor as in CLSE experiment. We then applied multi-class logistic regression for topic and emotion label classification. Lastly, we repeated the same with ground truth transcriptions provided by the authors of the database.
All three experiments were conducted in the 5-fold cross-validation setup on a subset of the dataset comprising 5,502 sentences labeled as ’angry’, ’happy’, ’sad’, and ’neutral’. The limited set of emotions is obtained as a standard pre-processing of IEMOCAP database.
VI IEMOCAP:Results
The results of our experiments and also results from comparable speech-based experiments taken from [15] are in Table IV.
@@ VI IEMOCAP:Results
These results not only highlight the robustness of textual features in emotion modeling but also invite further inquiry into the diminished role of paralinguistic elements. Intriguingly, our text based experiments yielded only $\sim 9\%$ absolutely worse results than the state-of-the-art large pre-trained models. This emphasizes the need for additional research to explore the complex interplay between textual and paralinguistic features in emotion classification.
@@ VII Discussion and conclusions
Reliance on ASR-focused pre-trained models like HuBERT risks conflating lexical and paralinguistic features. While proficient in text-dependent tasks, these models may obfuscate evaluations aimed at paralinguistic recognition, thereby exacerbating lexical overlap issues.
While our results suggest that the lexical features in the speech data significantly influence the classification metrics, they do not negate the importance of paralinguistic features altogether. In the tables, it is clear that when speech features are employed, indeed the performance of systems is better. They merely call for a more careful approach to evaluating paralinguistic recognition systems, with an explicit focus on decoupling textual and non-textual features.
We suggest that future work should focus on the development of methods for reducing text-dependency in existing datasets. Additional evaluations should also be conducted using datasets that have been explicitly designed to minimize lexical overlap, thereby offering a more reliable platform for paralinguistic studies.

### v2

@@ Abstract
Paralinguistic traits such as cognitive load and emotion are recognized as important areas in speech recognition research, often examined through specialized datasets such as CLSE and IEMOCAP. However, the integrity of these datasets is seldom scrutinized for text dependency. This paper critically evaluates the assumption that machine learning models trained on such datasets genuinely learn to identify paralinguistic traits, rather than merely capturing lexical features. By examining the lexical overlap in these datasets and testing the performance of machine learning models, we expose significant text dependency in trait labeling. Our results suggest that some machine learning models, especially large pre-trained models such as HuBERT, might inadvertently focus on lexical characteristics rather than the intended paralinguistic features.
@@ I Introduction
While the primary focus of speech recognition research gravitates towards Automatic Speech Recognition (ASR), the study of paralinguistic traits, such as cognitive load, physiological stress, and emotions, remains a significant field too. These traits are of interest for applications ranging from human-computer interaction to psychological research and rely heavily on dedicated datasets. Two key datasets are Cognitive Load with Speech and EGG (CLSE[1], commonly used for cognitive load recognition, and Interactive Emotional Dyadic Motion Capture (IEMOCAP[2], primarily employed for emotion recognition.
A prevailing assumption is that machine learning algorithms trained on these datasets learn to recognize paralinguistic traits based on observable physiological or psychological changes in speech production. However, this paper challenges this assumption by revealing a critical oversight: we provide evidence of significant lexical correlation between the labels (e.g., cognitive load or emotion) and the uttered sentences within these datasets.
We analyze CLSE and review the design of IEMOCAP to substantiate this claim. Given the implications of our findings, the paper serves as a call reevaluation of existing datasets and methodologies to ascertain that machine learning systems are learning to recognize what they are designed to recognize.
II Related Work
@@ II-A Datasets in Focus
  • CLSE (Cognitive Load with Speech and EGG): This is intended to test recognition of cognitive load by incorporating both speech recordings and Electroglottograph (EGG) signals.
@@
  • IEMOCAP (Interactive Emotional Dyadic Motion Capture) [^1] https://sail.usc.edu/iemocap/index.html: It covers about 12 hours of scripted and spontaneous dialogues. It captures speech and facial and hand movements to study a range of emotions.
II-B Machine Learning Approaches to Paralinguistics
@@
Various machine learning algorithms, from classic UBM-iVector [3], through LSTM [4] to recent large pre-trained models like wav2vec [5], have been employed on these datasets. The general presumption is that these models are capturing paralinguistic features rooted in physiological or psychological changes [6, 7].
Pepino et al [8] discovered the text-dependency issue in their work on Emotion Detection on IEMOCAP. In their experiments they observed over-optimistic results while using the original train-test splits. Yuanchao et al [9] analyzed wav2vec performance on IEMOCAP and found that this model prioritizes linguistic content over para-linguistic information.
@@
In paralinguistic research, it is essential to consider both the content and the manner of speech. Physiological and prosodic features often capture the manner of speech, while lexical choices, semantics, and syntax provide insights into the content. Ignoring either aspect may lead to an incomplete understanding of paralinguistic traits. This study emphasizes the importance of acknowledging lexical correlates when analyzing paralinguistic datasets to ensure comprehensive evaluation and accurate model training.
III Hypothesis and Methodology
III-A Hypothesis
Our hypothesis posits that machine learning models trained on paralinguistic speech recognition datasets like CLSE and IEMOCAP may exhibit significant text dependency. This dependency could undermine the models’ ability to genuinely learn paralinguistic traits, instead focusing on lexical characteristics of the speech data.
III-B Methodology
To test this hypothesis, we conducted experiments on both the CLSE and IEMOCAP datasets. We employed various machine learning models, including classic and state-of-the-art approaches, and analyzed their performance with and without lexical overlap. Our methodology involves steps seen in Figure 1.
AudioPreprocessing Whisper ASRSentence Embeddings LaBSEClustering AHCClassification Cluster IdentityEvaluation UARClassification MLREvaluation UARCLSEIEMOCAP
Figure 1: Block scheme of the experiment structure for different datasets
IV Dataset and Task Description
IV-A CLSE
@@
Our primary study focuses on the main part of the CLSE database, termed as CLSE-Span, which employs the  Reading Span task (RSPAN[10]. In RSPAN, participants are asked to validate sentences for their logical coherence while memorizing a sequence of letters. Sentences are randomly chosen from a closed set. The task is divided into sets, where each set contains varying trials of sentence validation and letter memorization. The database consists of 21 such sets with each participant producing 75 utterances, totaling 1800 utterances. The average utterance duration is four seconds. Labels for the cognitive load values are: low, medium, and high load.
IV-B IEMOCAP
The IEMOCAP dataset comprises 151 dialog videos, featuring a pair of speakers in each session, resulting in a total of 302 individual videos with approximately 12 hours of data. Ten actors were recorded in dyadic sessions (five sessions with two subjects each). They were asked to perform three selected scripts with clear emotional content. In addition to the scripts, the subjects were also asked to improvise dialogs in hypothetical scenarios, designed to elicit specific emotions (happiness, anger, sadness, frustration, and neutral state). The original annotations were expanded by the authors after the collection. The new annotated classes contain nine distinct emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral), along with metrics for valence, arousal, and dominance.
V Experimental Setup
V-1 Data Preprocessing
@@
We used Whisper ASR [11], a pre-trained multilingual large-v3 model, to transcribe the speech data from both datasets. Subsequently, we utilized Sentence Transformer embeddings as features, using the Language-Agnostic BERT Sentence Embeddings (LaBSE) model [12]. LaBSE utilizes a dual-encoder architecture trained in two steps: first on 17 billion monolingual sentences via Masked Language Modeling, and then further refined on six billion translation pairs covering 109 languages. The final model is publicly accessible [^2] https://huggingface.co/sentence-transformers/LaBSE and was used to extract 768-dimensional sentence embeddings.
V-2 Clustering
For the CLSE dataset, we applied Agglomerative Hierarchical Clustering (AHC) to the sentence embeddings to group similar sentences. After manually correcting minor clustering errors, we identified 81 different sentences. Each sentence was assigned an ID based on its cluster membership.
V-3 Classification
@@
Using the clustered data, we assigned the most frequent cognitive load label in the training set to each cluster in the CLSE dataset. These labels were used as a proxy to predict the cognitive load of utterances in the test and validation sets based on their cluster identity, effectively bypassing the use of any machine learning model for classification.
For the IEMOCAP dataset, we analyzed the recording sessions based on the type of scenario (improvised or scripted). We computed the most probable label per session, obtained ASR transcripts with Whisper ASR, and incorporated the sentence embeddings using the LaBSE model. We then applied multi-class logistic regression for topic and emotion label classification, and repeated the experiments using ground truth transcriptions provided by the dataset authors.
V-4 Evaluation
We evaluated the impact of lexical overlap by comparing performance metrics using the Unweighted Average Recall (UAR). This metric offers a balanced assessment of performance across all classes, making it suitable for evaluating imbalanced datasets. We compared the performance of models trained on the original dataset splits with those trained on fixed (shuffled) splits, where lexical redundancy was minimized.
V-A CLSE
We preprocessed, clustered and classified the CLSE dataset as described in previous section.
@@ V-A CLSE
V-B IEMOCAP
As the structure of the IEMOCAP database is different from CLSE, we have adopted a slightly modified approach to verify our hypothesis. IEMOCAP sessions are either improvised or scripted, with 12 different improvised scenarios and three scripts. We argue that the scenario of the recording session represents enough lexical context to the classifier that it can diminish the other factors (prosody etc.).
To verify this claim, we conceived three different experiments: In the first one, we computed the most probable label per given session in a similar fashion to Section V-A. Then we obtained ASR transcripts with Whisper ASR (large-v3) and incorporated the sentence embeddings using the same extractor as in the CLSE experiment. We then applied multi-class logistic regression for topic and emotion label classification. Lastly, we repeated the same with ground truth transcriptions provided with the database.
All three experiments were conducted in five-fold cross-validation setup on a subset of the dataset comprising 5,502 sentences labeled as ’angry’, ’happy’, ’sad’, and ’neutral’. The limited set of emotions is obtained as a standard pre-processing of IEMOCAP and corresponds to the author’s recommendations [13].
VI Results
VI-A CLSE
We assessed the influence of lexical overlap on classification performance using both original and fixed (shuffled) CLSE data splits. The original CLSE data displayed good performance metrics, but revealed patterns in the validation set that confirmed the lexical factors’ role.
@@ VI-A CLSE
After minimizing lexical redundancy in shuffled splits, the performance metrics notably decreased to the chance level, as can be seen in Table I, further substantiating the lexical overlap’s impact on classification.
We replicated the UBM-iVector system from the ComPaRE 2014 [14] challenge, using a 64-component Universal Background Model (UBM) with a 50-dimensional iVector extractor and Support Vector Machine (SVM) for classification [3]. Using the same approach as in the cluster-based approach, we obtained results showing significant degradation of the performance on validation and test sets (Table II).
@@
To corroborate our findings with a more recent speech-based system, we used HuBERT [15] (version large-ll60k) pre-trained on 60k hours from the Libri-Light dataset [16] as a feature extractor. As a classifier, we used an Attentive correlation pooling layer from [17]. This classifier uses a multi-head attention layer, where attention weights resemble a mixture model with heads parametrizing the mixture component. The results of our experiments are shown in Table III. They again manifest the same type of behavior as the previous experiments.
@@
VI-B IEMOCAP
For IEMOCAP, the results of our experiments and comparable speech-based experiments taken from [18] are shown in Table IV.
@@ VI-B IEMOCAP
[…]
@@ VII Discussion and Conclusions
Reliance on ASR-focused pre-trained models like HuBERT risks conflating lexical and paralinguistic features. While excelling in text-dependent tasks, these models may obfuscate evaluations aimed at paralinguistic recognition.
While our results suggest that the lexical features in the speech data significantly influence the classification metrics, they do not negate the importance of paralinguistic features altogether. It is clear that when speech features are employed, indeed the performance of systems is better. They merely call for a more careful approach to evaluating paralinguistic recognition systems, with an explicit focus on decoupling textual and non-textual features.
We suggest that future work should focus on the development of methods for reducing text dependency in existing datasets. Additional evaluations should also be conducted using datasets that have been explicitly designed to minimize lexical overlap.